# Using an LLM to Turn Sign Spottings into Spoken Language Sentences

### v1

@@ IV-D Qualitative results
Morgen fahre ich mit dem Fahrrad zur Uni, weil mein Dozent sagt, frische Luft ist wichtig. (Tomorrow I’ll ride my bike to university because my lecturer says fresh air is important.)
V CONCLUSIONS AND FUTURE WORKS
In this paper, we proposed a sign language translation method integrating a sign spotter with the natural language generation capabilities of ChatGPT. Our approach does not need any training or fine-tuning for sign language translation. Also, it is capable of producing coherent and fluid sign language sentences.
@@ V CONCLUSIONS AND FUTURE WORKS
One of the limitations of our approach is that the performance is dependent on the vocabulary of our spotter. However, our approach can be adapted to specialized sign language interpretation by fine-tuning the spotter on domain-specific gloss sets. A future direction may include expanding the vocabulary of the spotter to increase the range of recognized glosses. By increasing the range of recognized glosses, our system can become even more versatile and applicable across various sign language contexts.
VI ACKNOWLEDGMENTS
This work was supported by the SNSF project ‘SMILE II’ (CRSII5 193686), European Union’s Horizon2020 programme (‘EASIER’ grant agreement 101016982) and the Innosuisse IICT Flagship (PFFS-21-47). This work reflects only the authors view and the Commission is not responsible for any use that may be made of the information it contains. All experiments were run, datasets accessed, and models built by researchers at the University of Surrey.

### v3

@@ 4.3. Quantitative Results
Spotter+GPT Morgen fahre ich mit dem Fahrrad zur Uni, weil mein Dozent sagt, frische Luft ist wichtig. (Tomorrow I’ll ride my bike to university because my lecturer says fresh air is important.)
4.4. Qualitative results
We provide qualitative results of our approach in Table 5. When the spotter successfully detects the majority of glosses, GPT effectively generates high-quality spoken sentences from the glosses (Example 1, 2, and 4). Even in cases where the glosses are incomplete or contain additional glosses, ChatGPT still preserves semantic coherence with information gaps or the generation of new content (Examples 5, and 6). On the other hand, not surprisingly, ChatGPT’s performance heavily depends on the quality of the input glosses. When the spotter fails to detect glosses, ChatGPT generates incorrect sentences (Example 3) or “No translation”.
5. Conclusion
In this paper, we proposed a novel sign language translation framework that combines a sign spotter with a Large Language Model (LLM), specifically ChatGPT, to generate spoken language sentences from sign language videos. Our method does not require any end-to-end SLT model training and leverages prompt-based inference for gloss-to-text generation.
Experimental results on MeineDGS-V and a newly collected DGS-20 dataset show that Spotter+GPT produces coherent and semantically accurate sentences, especially when the spotter successfully identifies relevant glosses. This indicate that leveraging LLMs offers a promising and flexible alternative to traditional gloss-to-text pipelines.
Our system can process video inputs from both pre-recorded datasets and live capture devices such as webcams. This flexibility allows us to support potential real-time applications.
@@ 5. Conclusion
However, our system’s performance is inherently constrained by the vocabulary and accuracy of the sign spotter. When the spotter fails to detect critical glosses, the translation quality drops significantly. This highlights the importance of high-quality gloss spotting. Nevertheless, our approach can be adapted to specialized sign language interpretation by fine-tuning the spotter on domain-specific gloss sets. A future direction may include expanding the vocabulary of the spotter to increase the range of recognized glosses.
Acknowledgements.
We would like to thank Necati Cihan Camgoz for the valuable discussions and feedback. This work was supported by the SNSF project ‘SMILE II’ (CRSII5 193686), the Innosuisse IICT Flagship (PFFS-21-47), EPSRC grant APP24554 (SignGPT-EP/Z535370/1), and through funding from Google.org via the AI for Global Goals scheme. This work reflects only the author’s views and the funders are not responsible for any use that may be made of the information it contains.